# One Framework to Rule Them All: Unifying Multimodal Tasks with LLM Neural-Tuning

### v2

@@ 4 MMUD Benchmark
We constructed our dataset using open-source datasets: RefCOCO [27], RefCOCOg [28], and RefClef [29, 28]. Since these datasets only provide referring segmentation masks, we augmented them with annotations for other tasks. The dataset generation pipeline is illustrated in Figure 3. Initially, we employed GPT-4v to generate image captions describing the contents. Subsequently, GPT-4v was used to generate complex questions along with answers based on the image contents. We then performed manual filtration to ensure that the generated contents were meaningful and suitable for multimodal understanding. We filtered out samples in cases where: 1) the generated contents were meaningless; 2) the length of the generated contents was either too long or too short; 3) the answers could be directly inferred from the questions without reference to the images. To enhance the capability of complex reasoning, we manually inserted the $<$OBJ-i$>$ token after each object in the answer, where $i$ represents the $i$-th object in the image. For example, when generating the question Can you judge the relationship between the two people in the front based on what happened in the picture?” with the answer with answer In this image, one of the women in the front is cutting a pizza, and a little girl is next to her…, we manually inserted the $<$OBJ$>$ token after the expression of each object: In this image, one of the women in the front $<$OBJ-1$>$ is cutting a pizza $<$OBJ-2$>$, and a little girl $<$OBJ-3$>$ is next to her…. Following the structure of the RefCOCO datasets, each sample image contains multiple meaningful objects with corresponding masks. Each $<$OBJ-i$>$ token not only corresponds to an object but also to the corresponding segmentation mask, which could help models better capture the relationship between texts and images. The dataset will be made publicly available to advance research in this area.
@@ 5.2 Results on MMUD
Table 1 and Table 2 presents the results of our proposed neural tuning compared with previous methods. For vanilla referring segmentation and reasoning segmentation, we mainly compare our method with LVAT, LISA, and PixelLM. We employ the mean intersection over union (mIoU) and overall IoU as the metrics. LVAT aims to fuse the BERT features into vision backbones, achieving great results in vanilla segmentation. However, when it comes to complex reasoning scenarios, LVAT fails to converge while other LLM-based methods yield good results, demonstrating the powerful reasoning capabilities of large models. Compared with LISA and PixelLM, which employ pretrained LLaVA [12] and focus on complex reasoning segmentation, neural tuning can achieve better results, demonstrating its effectiveness. Regarding inference speed, our tuned 7B model can complete the reasoning segmentation process within 80ms per text-image pair on an RTX 4090 GPU, while the 13B model takes 110ms on an RTX A6000 GPU. This demonstrates the efficiency of our proposed neural tuning. We believe this efficiency results from the linear sparse vector updating and task guidance.
Furthermore, to further evaluate the effectiveness of our proposed method, we also present the performance on the original test set of RefCOCO, RefCOCO+, and RefCOCOg datasets  [^2] For fair comparision, we re-split the train, validation, and test set of MMUD according to the original datasets.. As shown in Table 1(b), our method achieves state-of-the-art performance on the public datasets but with lower computational burden, revealing the effectiveness and efficiency of our proposed multitask neural tuning for segmentation tasks.
@@
On text-to-image synthesis tasks, quantitative metrics are shown in Table 2. We employ FID score [35], KID Score [36], and inception score (IS [37]) as the metrics. We compare our method with previous methods, including Parti [38], GLIGEN [39], and U-ViT-S/2 [40]. Compared with other methods, one of the significant advantage of our method is that neural tuning is designed for multitask tuning instead of a certain task.
@@ 5.5 Ablation Study
Choice of image generation decoders: Apart from VQGAN, there are alternative options for image decoding. We also utilized a pre-trained DALL-E2 decoder [^3] http://github.com/hidden/for/review for generating images from the hidden embeddings. A comparison is presented in Figure 4 (Section  5.3). The quality and style of the generated images are significantly influenced by the pre-trained decoders. This bias is a result of the pretraining data used for the decoders. Hence, in theory, neural tuning can employ any pre-trained image synthesis decoder. Unfortunately, due to hardware limitations, we were unable to test other pre-trained models for image generation.
Table 4: The ablation of $\beta$ in sparse vector updating.
Tasks involved oIoU of Reasoning Segmentation on MMUD
@@
Ablation of $\beta$ in Sparse Vector Updating The detailed ablations regarding $\beta$ in sparse vector updating are presented in Table 4. In these experiments, we consistently use the reasoning segmentation task and employ the corresponding oIoU for evaluation. For instance, two tasks involved refer to reasoning segmentation and vanilla segmentation, while only one task involved refers to reasoning segmentation only. From the results, we observe that setting $\beta$ to 0.4 yields the best performance when all four tasks are involved. For scenarios involving three tasks, the optimal performance is achieved with $\beta$ set to 0.5 or 0.6. Finally, when there are only one or two tasks, it is more beneficial to activate a higher proportion of neurons, with $\beta$ set to 0.9 or 1.0.
@@ 6 Conclusion
In this paper, we present a novel unified multimodal multitask learning framework with a new tuning strategy called neural tuning. Under this framework, we unify tasks using an all-in-token approach, which enhances scalability by facilitating the integration of additional modalities or tasks. Additionally, we emulate human cognitive processes through sparse distributed representation, activating specific neurons for different tasks. We evaluate our method across four tasks, including reasoning segmentation and text-to-image synthesis, and demonstrate competitive performance compared to current state-of-the-art methods. To support further research in this area, we introduce the MMUD dataset, which provides a diverse set of annotations for various tasks. Moving forward, we plan to explore additional modalities, such as acoustics, and expand our research into fields like medical applications.
Appendix A Example Appendix Section
Appendix text.

### v3

@@ 4 MMUD Benchmark
We constructed our dataset using open-source datasets: RefCOCO [22], RefCOCOg [23], and RefClef [24, 23]. Since these datasets only provide referring segmentation masks, we augmented them with annotations for other tasks. The dataset generation pipeline is illustrated in Figure 4. Initially, we employed GPT-4v to generate image captions describing the contents. Subsequently, GPT-4v was used to generate complex questions along with answers based on the image contents. We then performed manual filtration to ensure that the generated contents were meaningful and suitable for multimodal understanding. We filtered out samples in cases where: 1) the generated contents were meaningless; 2) the length of the generated contents was either too long or too short; 3) the answers could be directly inferred from the questions without reference to the images. To enhance the capability of complex reasoning, we manually inserted the <<OBJ-i>> token after each object in the answer, where $i$ represents the $i$-th object in the image. For example, when generating the question Can you judge the relationship between the two people in the front based on what happened in the picture?” with the answer with answer In this image, one of the women in the front is cutting a pizza, and a little girl is next to her…, we manually inserted the <<OBJ>> token after the expression of each object: In this image, one of the women in the front <<OBJ-1>> is cutting a pizza <<OBJ-2>>, and a little girl <<OBJ-3>> is next to her…. Following the structure of the RefCOCO datasets, each sample image contains multiple meaningful objects with corresponding masks. Each <<OBJ-i>> token not only corresponds to an object but also to the corresponding segmentation mask, which could help models better capture the relationship between texts and images. The dataset is publicly available at https://github.com/kiva12138/NeuralTuning.
@@ 5.1 Experimental Settings
Table 1: The summary of our our evaluated tasks, datasets, metrics, and corresponding results.
Tasks Datasets Main Metric Main Results
Referring
Segmentation
RefCOCO
oIoU
74.2
RefCOCO+ 66.3
RefCOCOg 70.3
Reasoning
Segmentation
MMUD 64.2
Image
Captioning
BLEU-4 43.7
Image
Synthesis
FID 12.7
VQA
VQA v2
VQA
72.1
TextVQA 57.6
@@ 5.2 Results on MMUD
Table 1 summarizes the evaluated tasks, datasets, main metric and corresponding results for our proposed framework. In detail, Table 2, Table 3, and Table 4 presents the comparison with previous methods. For vanilla referring segmentation and reasoning segmentation, we mainly compare our method with LAVT [12], LISA [14], and PixelLM [2]. We employ the mean intersection over union (mIoU) and overall IoU as the metrics. LAVT aims to fuse the BERT features into vision backbones, achieving great results in vanilla segmentation. However, when it comes to complex reasoning scenarios, LAVT fails to converge while other LLM-based methods yield good results, demonstrating the powerful reasoning capabilities of large models. Compared with LISA and PixelLM, which employ pretrained LLaVA [9] and focus on complex reasoning segmentation, neural tuning can achieve better results, demonstrating its effectiveness. Regarding inference speed, our tuned 7B model can complete the reasoning segmentation process within 80ms per text-image pair on an RTX 4090 GPU, while the 13B model takes 110ms on an RTX A6000 GPU. This demonstrates the efficiency of our proposed neural tuning. We believe this efficiency results from the linear sparse vector updating and task guidance.
Furthermore, to further evaluate the effectiveness of our proposed method, we also present the performance on the original test set of RefCOCO, RefCOCO+, and RefCOCOg datasets  [^2] For fair comparison, we re-split the train, validation, and test set of MMUD according to the original datasets.. As shown in Table 2(b), our method achieves state-of-the-art performance on the public datasets but with lower computational burden, revealing the effectiveness and efficiency of our proposed multitask neural tuning for segmentation tasks.
@@
Table 4: The zero-shot performance of our proposed framework with previous methods on VQA v2 (test-dev) and TextVQA (val) datasets for visual question answering task.
Method VQA v2 TextVQA
BLIP-2 [7] 41.0 42.5
InternVL-Chat [25] 72.3 42.1
InstructBLIP [26] - 50.1
SPHINX-Intern2 [27] 75.5 58.1
VisionLLM v2 [11] 80.8 64.7
NT-7B(Ours) 72.1 57.6
@@
On text-to-image synthesis tasks, quantitative metrics are shown in Table 3. We employ FID score, KID Score, and inception score (IS) as the metrics. We compare our method with previous methods, including Parti [33], GLIGEN [34], and U-ViT-S/2 [35]. Compared with other methods, one of the significant advantage of our method is that neural tuning is designed for multitask tuning instead of a certain task.
In addition, we evaluated our proposed method on visual question answering (VQA) tasks in a zero-shot manner. The experiments are conducted on VQA v2 [36] and TextVQA [37] datasets. The results, presented in Table 4, demonstrate the effectiveness of our approach in handling visual question answering tasks in a zero-shot manner, without task-specific fine-tuning. Notably, our method achieves competitive performance compared to state-of-the-art models such as VisionLLM v2 [11] and InterVL-Chat [25], despite not relying on specialized tuning for individual tasks. This highlights the strong generalizability and robustness of our framework across diverse benchmarks, emphasizing its practicality for multitask and low-resource scenarios.
Despite the promising results demonstrated by our unified framework, particularly in segmentation tasks where it achieves state-of-the-art performance, we observe trade-offs in its performance on generative tasks such as image synthesis. While the framework matches or slightly underperforms compared to specialized methods [33, 34] in these tasks, this outcome reflects the inherent balance between generality and specialization. Unlike task-specific architectures optimized exclusively for image generation [35], our unified framework prioritizes efficiency, flexibility, and scalability across diverse tasks. This design choice ensures robust multitask learning capabilities but may limit its ability to fully exploit the potential of generative tasks. Future work could address this trade-off by integrating advanced generative modules into the framework, enabling enhanced performance while preserving the lightweight and unified structure.
@@ 5.5 Ablation Study
Choice of image generation decoders: Apart from VQGAN, there are alternative options for image decoding. We also utilized a pre-trained DALL-E2 decoder for generating images from the hidden embeddings. A comparison is presented in Figure 5 (Section  5.3). The quality and style of the generated images are significantly influenced by the pre-trained decoders. This bias is a result of the pretraining data used for the decoders. Hence, in theory, neural tuning can employ any pre-trained image synthesis decoder. Unfortunately, due to hardware limitations, we were unable to test other pre-trained models for image generation.
Table 6: Ablation study of $\beta$ and $\alpha$ in sparse vector updating. The metrics reported are the oIoU scores for Reasoning Segmentation on the MMUD dataset.
Tasks involved Ablation on the Selection of $\beta$
@@
4 Task Ablation on the Selection of $\alpha$
$\alpha=0.1$ $\alpha=0.2$ $\alpha=0.3$ $\alpha=0.4$ $\alpha=0.5$ $\alpha=0.6$ $\alpha=0.7$ $\alpha=0.8$ $\alpha=0.9$ $\alpha=1.0$
55.7 57.6 59.2 61.1 59.5 60.2 61.1 61.0 62.8 59.3
Ablation of $\beta$ and $\alpha$ in Sparse Vector Updating The detailed ablations regarding $\beta$ and $\alpha$ in sparse vector updating are presented in Table 6. In these experiments, we consistently use the reasoning segmentation task and employ the corresponding oIoU for evaluation. For instance, two tasks involved refer to reasoning segmentation and vanilla segmentation, while only one task involved refers to reasoning segmentation only. From the results, we observe that setting $\beta$ to 0.4 yields the best performance when all four tasks are involved. For scenarios involving three tasks, the optimal performance is achieved with $\beta$ set to 0.5 or 0.6. Finally, when there are only one or two tasks, it is more beneficial to activate a higher proportion of neurons, with $\beta$ set to 0.9 or 1.0. For the selection of $\alpha$ in EMA, we find that performance is optimal when $\alpha$ is set to 0.9, suggesting that cross-layer updating is more stable at this value. Conversely, smaller $\alpha$ values result in a dramatic performance drop, likely due to numerical instability in the cross-layer guidance updates. As a result, we set $\alpha$ to 0.9 in guidance signal updating.
@@ 6 Conclusion
In this paper, we present a novel unified multimodal multitask learning framework with a new tuning strategy called neural tuning. Under this framework, we unify tasks using an all-in-token approach, which enhances scalability by facilitating the integration of additional modalities or tasks. Additionally, we emulate human cognitive processes through sparse distributed representation, activating specific neurons for different tasks. We evaluate our method across four tasks, including reasoning, segmentation, and text-to-image synthesis, demonstrating competitive performance compared to current state-of-the-art methods. To support further research in this area, we introduce the MMUD dataset, which provides a diverse set of annotations for various tasks. The tuned model weights will also be made publicly available to foster innovation and collaboration in this domain.
6.1 Current Limitations
Despite the significant contributions of our work, certain limitations remain. Firstly, the current implementation does not incorporate acoustic modalities, which are critical for applications such as speech analysis and audio-visual reasoning. Secondly, while our framework has demonstrated strong performance across four tasks, its potential scalability to a broader range of tasks, including more than 10 simultaneous tasks, remains unexplored. The computational cost for such large-scale multitask training may become a concern, particularly for researchers with limited resources.
Furthermore, our framework currently employs LLaMA2 as the baseline LLM for experiments, leveraging both the 7B and 13B variants. While LLaMA2 is a strong and widely used model, it is not the most recent state-of-the-art in large language models. We acknowledge that adopting newer and more advanced LLMs could further enhance performance across tasks, particularly in scenarios demanding greater reasoning and generative capabilities. However, our experiments on multiple scales of LLaMA2 demonstrate that larger models consistently improve results, underscoring the scalability and generalizability of our approach.
Additionally, the potential effects of task token overlap or interference were not a major concern in our experiments with four tasks using 128-dimensional sparse vectors. However, as the number of tasks increases, higher-dimensional sparse vectors may be required to mitigate interference, representing a theoretical aspect that warrants further investigation. Furthermore, the dependency on decoder choice for tasks like image synthesis was only partially explored in this work, as we evaluated VQGAN and DALL-E2 due to time constraints. Exploring alternative decoders could reveal additional insights into the generalizability of our framework.
6.2 Future Research
Our contributions lay a robust foundation for future advancements in multimodal multitask learning. The proposed MMUD dataset provides a valuable benchmark for evaluating such frameworks, and the release of model weights ensures reproducibility and facilitates adaptation across diverse applications. By leveraging a cognitive-inspired sparse task network and emphasizing parameter efficiency, our work highlights the intersection of neuroscience principles and machine learning methodologies.
Looking ahead, we aim to address the aforementioned limitations. This includes integrating acoustic modalities, extending the framework to accommodate tasks across diverse domains such as medical diagnostics, autonomous systems, and creative content generation, and conducting more comprehensive evaluations on public benchmarks. We also plan to optimize the framework to enhance its computational efficiency, ensuring accessibility for a wider range of researchers. Finally, we will also explore integrating the framework with cutting-edge LLMs, such as QWEN 3 or other advanced models, to fully exploit the latest advancements in multimodal understanding and multitask learning. These efforts will strengthen the robustness, scalability, and impact of our framework, further advancing the frontiers of multimodal multitask learning.
Appendix A Data Filtration and Annotation Guidelines
To ensure the quality and suitability of the MMUD dataset for multimodal understanding, we employed a rigorous data filtration and annotation process following the initial data generation via GPT-4. Below, we outline the steps taken to refine the dataset and address potential issues in the generated content.
A.1 Data Filtration Process
After generating initial annotations using GPT-4, we manually reviewed and filtered the samples based on the following criteria:
  • 1.
    Meaningfulness: Samples were excluded if the generated content was nonsensical or lacked coherence (for example: a simple response Yes, I know that answer.).
  • 2.
    Length Appropriateness: Samples were filtered out if the length of the generated content was excessively long (longer than 500 words) or too short (shorter than 10 words) to provide meaningful multimodal understanding.
  • 3.
    Image Reference Dependency: Samples were discarded if the answers could be directly inferred from the questions without requiring reference to the associated images.
A.2 Enhancement for Complex Reasoning
To improve the dataset’s capability to support complex reasoning tasks, we introduced object tokens (<<OBJ-i>>) in the generated answers. Each <<OBJ-i>> token corresponds to a specific object in the image, enabling the models to better align textual descriptions with visual content.
For example, for a question such as:
Can you judge the relationship between the two people in the front based on what happened in the picture?
The initial answer was:
In this image, one of the women in the front is cutting a pizza, and a little girl is next to her…
We manually inserted object tokens to provide a more structured and precise representation:
In this image, one of the women in the front <<OBJ-1>> is cutting a pizza <<OBJ-2>>, and a little girl <<OBJ-3>> is next to her…
(a) The caption length distribution in MMUD.
(b) The reasoning QA pair length distribution in MMUD.
(c) The object counts per sample in MMUD.
Figure 7: The distribution of caption length, reasoning question-answer pair length, and object counts for each sample in MMUD.
Table 8: The comparison of neural tuning, LoRA, and prefix tuning on RefCOCO, RefCOCO+, and RefCOCOg. The metric used in the table is oIoU.
Method TFLOPs RefCOCO RefCOCO+ RefCOCOg(U)
Val TestA TestB Valid TestA TestB Valid Test
Prefix Tuning-7B 3.32 70.3 72.7 65.0 61.1 67.8 55.7 67.1 66.2
LoRA-7B 2.38 72.1 74.2 67.3 64.2 69.9 56.4 68.0 67.1
NT-7B(Ours) 2.47 74.2 76.7 68.0 66.3 71.2 58.1 70.3 70.2
A.3 Object Annotations and Masks
To ensure the quality and utility of the MMUD dataset, we manually annotated all 36,582 samples, with each sample containing descriptions of 2.8 objects in average. For each object, we inserted an <<OBJ-i>> tag after its corresponding description to facilitate multimodal reasoning and segmentation tasks. During the process, we ensured that each tag corresponds to its respective mask in the RefCOCO series datasets. This annotation process was completed over a period of two months by three dedicated annotators. While labor-intensive, this effort was crucial for creating a high-quality dataset that aligns with the goals of multimodal multitask learning research. The detailed annotations provide a robust foundation for various tasks, ensuring that the dataset is both comprehensive and reproducible for the research community.
After data annotation, to provide a comprehensive overview of the MMUD dataset, we present key metrics that highlight its diversity and richness. Figure 7(a) illustrates the distribution of caption lengths, showing that most captions fall within the range of 50 to 200 tokens, reflecting the dataset’s emphasis on detailed and descriptive annotations. Similarly, Figure 7(b) depicts the length distribution of reasoning question-answer pairs, with the majority spanning 100 to 400 tokens, underscoring the complexity and depth of reasoning required for these tasks. Lastly, as shown in Figure 7(c), the object counts per sample reveal that most samples contain fewer than three objects, with two objects being the most common. This distribution aligns with the design of the dataset to balance between simplicity and complexity, ensuring practical usability across a range of multimodal applications.
Appendix B Comparison of Neural Tuning with LoRA and Prefix Tuning
We evaluated the parameter efficiency and performance of our neural tuning mechanism against common alternatives, including Low-Rank Adaptation (LoRA) [38] and prefix tuning [40]. These experiments on vanilla referring segmentation tasks. The results are summarized in Table 8. Neural tuning demonstrates similar TFLOPs to LoRA, indicating comparable computational efficiency, while consistently achieving superior performance across all evaluated metrics. Specifically, neural tuning achieves an oIoU of 74.2%, compared to 72.1% for LoRA and 70.3% for prefix tuning.
These results highlight neural tuning as a balanced and scalable solution for multitask and multimodal learning. Its advantage stems from dynamically allocating computation while maintaining lightweight operations, surpassing traditional tuning strategies in both performance and efficiency.
Table 9: The multitask interference analysis for tasks involved when activating all neurons in sparse vectors.
Main Task Additional Task Performance $\Delta$
Referring Segmentation (oIoU on RefCOCO TestA Set) - 75.4 -
Reasoning Segmentation 76.0 +0.6%
Imgae Captioning 75.4 +0.0%
Text-to-Image Generation 74.9 -1.5%
Reasoning Segmentation (oIoU on MMUD Test Set) - 61.7 -
Referring Segmentation 62.0 +0.3%
Imgae Captioning 61.1 -0.6%
Text-to-Image Generation 60.4 -1.3%
Imgae Captioning (BLEU-4 on MMUD Test Set) - 44.1 -
Reasoning Segmentation 44.0 -0.1
Referring Segmentation 49.7 -0.4
Text-to-Image Generation 42.8 -1.3
Text-to-Image Generation (FID($\uparrow$) on MMUD Test Set) - 11.9 -
Reasoning Segmentation 12.8 +0.9
Referring Segmentation 13.0 +1.1
Image Captioning 12.2 +0.3
Appendix C Multitask Interference Analysis
This section presents experiments analyzing task interference when all neurons are activated during multitask training. The objective was to examine how tasks interact within a unified framework.
We evaluated four tasks: referring segmentation, reasoning segmentation, text-to-image generation, and captioning, under full neuron activation without sparsity. As shown in Table 9, semantically related tasks—such as referring and reasoning segmentation—benefited from shared features like spatial reasoning and object localization, yielding oIoU gains of 0.6% and 0.3%, respectively. In contrast, tasks with divergent goals exhibited negative interference. Reasoning segmentation and text-to-image generation showed conflicting activation patterns, resulting in a 0.3% drop in oIoU and a 0.3 increase in FID, respectively. Captioning, relying on more distinct neuron subsets, showed minimal interaction, though slight performance fluctuations suggest mild cross-task interference.
To address task interference, we adopted a neuroscience-inspired sparse neuron activation strategy that selectively activates neurons per task. This approach mitigates negative interactions while preserving beneficial correlations, leading to stable performance across tasks. The sparsity parameter $\beta$, empirically set to 0.4 (see Section 5 and Table 6), effectively reduces overlap and supports consistent, efficient multitask learning.
These findings underscore the value of task-specific strategies in multitask frameworks. By minimizing neuron overlap, our neural tuning method strikes a balance between task isolation and shared learning, boosting both individual and overall performance.